OpenStudy (anonymous):

How do you convert the angle from radians to degrees? How do you find the complement of the angle? And How do you find the supplement of the angle???

OpenStudy (anonymous):

angle from radians to degrees ----> multiply the angle by 180/pi compliment of angle = 90- angle (compliment of 50 =90-50=40) supplement of angle = 180- angle (supplement of 100=180-100 =80)

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

any specific angles ?

OpenStudy (anonymous):

yea hold on let me find them

OpenStudy (anonymous):

this is convert the angle from radians to degrees 5pi over 6 and then 4pi over 3

OpenStudy (anonymous):

then find th complement of the angle is pi over 6

OpenStudy (anonymous):

and find the supplement of the angle 161 degrees 18'49"

OpenStudy (anonymous):

5pi / 6 * 180/pi =.... ? 4pi / 3 *180/pi =... ?

OpenStudy (anonymous):

5pi / 6 * 180/pi = 5*180/6 = 150 degrees try for 4pi/3

OpenStudy (anonymous):

150?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

now for 4pi/3 ?

OpenStudy (anonymous):

2368.705056?

OpenStudy (anonymous):

omg, what have you done :O

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

walk me though it please

OpenStudy (anonymous):

4pi/3 = 4pi/3 * 180/pi = 4*180/3 = ... ?

OpenStudy (anonymous):

240?

OpenStudy (anonymous):

to convert to degrees, you multiply the angle by 180/pi

OpenStudy (anonymous):

okay i got you but is 240 right?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

yay!! now can you help with the others

OpenStudy (anonymous):

sure complement of the angle is pi over 6 complimentary angles add upto 90 degrees or pi/2 radians so, compliment of pi/6 will be pi/2- pi/6 =.... ?

OpenStudy (anonymous):

pi/2 - pi/6?

OpenStudy (anonymous):

would it be 30 degrees?

OpenStudy (anonymous):

do you need answer in degrees ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

pi/2-pi/6 = 90- 30 = 60 degrees

OpenStudy (anonymous):

so, compliment of pi/6(30 degrees) is 60 degrees

OpenStudy (anonymous):

okay thank you and the last one..and i think i wanna see if an answer is right with you if thats okay

OpenStudy (anonymous):

its ok with me.

OpenStudy (anonymous):

you still there?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

supplement of the angle 161 degrees 18'49" to get supplement, just subtract the angle from 180 degrees 180-161 18'49" = .... ?

OpenStudy (anonymous):

can you explain a little bit more please

OpenStudy (anonymous):

would it be 19.31361....

OpenStudy (anonymous):

like supplementary angles add upto 180 degrees so, if i have 100 degrees,its supplement is 80 right ? because 100+80 = 180 how to find that ? , just subtract the angle from 180

OpenStudy (anonymous):

so would it be 19.31361.....?

OpenStudy (anonymous):

the answer should be like 180 degrees 0' 0"- 161 degrees 18' 49" = 18 degrees 42' 11" i got 11 and 42 by subtracting 49 and 18 from 60

OpenStudy (anonymous):

okay so the answer would be 18 degrees 42' 11"

OpenStudy (anonymous):

would that be right?? cause it says find the supplement of the angle 161degrees 18'49"

OpenStudy (anonymous):

yeah